1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

The Banshee (An Approach)

The Werewolf El Hechice- ro

Varulven

tradukita de Max Knight tradukita de Alexander Gross tradukita de Hans Eiseneck tradukita de Sven Collberg

One night, a banshee

slunk away from mate and child,

and in the gloom went to a village tea-

cher’s tomb, requesting him: “Inflect

me, pray.”

A Werewolf, troubled by

his name, Left wife and brood one

night and came To a hidden graveyard to

enlist The aid of a long-dead phi-

lologist.

Un hechicero, con

esposa vino un día a la fo-

sa de su maestro, y al

profesor pidió: “¿Me cuen-

tas? ¡Por favor!”

En varulv lopp från fru m.m. en vacker afton samt

begav sig till en bymagisters grav och bad: “Var snäll och konjugera.

The village teacher

climbed up straight upon his grave stone

with its plate and to the apparition

said who meekly knelt befo-

re the dead:

“Oh sage, wake up, please

don’t berate me,” He howled sadly, “Just con-

jugate me.” The seer arose a bit unstea-

dy Yawned twice, wheezed on-

ce, and then was ready.

El profesor sin va-

cilar subió y comenzó a

hablar. Desde allá de su

letrero se dirigió al hechi-

cero: mig!” Byskolläraren stod upp på blecknamnsskyltens mässingknopp med svar till den, som satt på pass med tåligt korslagd tass på tass.

“The banSHEE, in the

subject’s place; the banHERS, the pos-

sessive case. The banHER, next, is

what they call objective case–and that

is all.”

“Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singu-

larly cast: There’s ‘Amwolf’ too, the

present tense, And ‘Iswolf,’ ‘Arewolf’ in

this same sense.”

“El hechiuno” le

explica, “el hechidos” de-

spués indica, “el hechitres” dice

al fin, “con eso basta

¡chiquitín!”

“Jag var-ulv 1:a pers.

i sing., du, han, hon, den, det var-ulv. Så vi voro-ulv, I voren.

På de voro-ulv kom ingenting.”

The banshee marveled

at the cases and writhed with plea-

sure, making faces, but said: “You did not

add, so far, the plural to the singu-

lar!”

“I know that–I’m no mental

cripple– The future form and parti-

ciple Are what I crave,” the be-

ast replied. The scholar paused–again

he tried:

El hechicero, con-

movido por lo que capta

su oído, al maestro ruega,

encantado, que le adjunte un

quebrado.

Varulven myste glad

åt hur pers. hade följts av sing. och plur. “Fast”, bad han, “kan det inte ges ens en enda pers. i tempus presens?”

…  …  …  …

2/3

The teacher, though,

admitted then that this was not within

his ken. “While bans are fre-

quent”, he advised, “A she cannot be pluri-

zed.”

“A ‘Will-be-wolf?’ It’s just

too long: ‘Shall-be-wolf?’ ‘Has-been-

wolf?’ Utterly wrong! Such words are wounds

beyond all suture– I’m sorry, but you have no

future.”

Pero el profesor

confiesa que nada sabe con

certeza: “Hay muchos he-

chos. Hechiceros

únicamente hay enteros.”

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine im-

perfecto var contradictio in

adjecto.

The banshee, rising

clammily, wailed: “What about

my family?” Then, being not a lear-

ned creature, said humbly “Thanks”

and left the teacher.

The Werewolf knew better–

his sons still slept At home, and homewards

now he crept, Happy, humble, without

apology For such folly of philology.

El brujo se volvió

muy triste: “Mi cara mitad,

¿no la viste?” Ya que no era sa-

bio con gratitud se de-

spidió.

Varulven satt med tårad lins. Han visste att han fanns och finns. Därtill kom också fru

m.m. som även måste exi-

stera.

…  …  …  …

3/3

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en hispana de Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .